# Active Light Modulation to Counter Manipulation of Speech Visual Content

### v1

@@ 3. Preliminaries
Spotlight is a physical and proactive approach to protecting live speech videos from visual falsification of speaker identity as well as lip and face movements, which reflect delivered content and speech speed. To match the visual nature of these falsifications, we embed physical signatures with light – the sensing modality of vision. In this section, we describe Spotlight’s design and threat model, and then discuss the technical challenges it overcomes.
@@ 3.1. System Overview
Figure 3 overviews the Spotlight design, comprising three modules. The Spotlight core unit, deployed at the site of a live speech, consists of a camera observing the event and light modulator positioned to project light onto any approximately planar surface in the immediate vicinity of the speaker (e.g., a small portion of a podium, wall, backdrop, or curtain). Notably, each of the speech environments portrayed in recent high-profile falsification incidents (nancy19, ; hadid23, ; swinney2024, ; biden23_draft, ; biden24_lipsync, ; harris24_slow, ; harris24_slur, ; pelosi_trump_tweet, ) possess such a surface. We believe that, in practice, it is rare to find a high-profile speech event where this is not the case.
For each window of camera video frames, the signature creation module running on the core unit extracts a digest from frames. The digest contains semantically-meaningful features capturing speaker identity and lip and face motion and additional provenance metadata such as a window timestamp. The core unit then generates a message authentication code (MAC) for the digest with its secret key. A digest and its MAC comprise a signature. The optical signature embedding module encodes the signature data as optical modulations projected into the scene in the subsequent window of time. Thus, signatures are naturally embedded into real recordings.
To verify a published video of a speech, the verification module extracts optical signatures from video frames and validates the MACs to confirm digests’ integrity. It then examines whether the semantically-meaningful features recovered from the digests match those of the portrayed speech. To ensure that the secret key used to validate MACs is secured, we envision the verification module as a secure cloud service queried by users and media platforms. As such, secret keys will never be exposed to third-party devices. We elaborate on Spotlight’s use of secret keys and MACs in §4.3.
@@ 3.2. Threat Model
Video producers create and disseminate legitimate videos of the speech. This group includes viewers recording at the event and non-malicious parties re-distributing videos. Audience members can record using any RGB camera with a frame rate $\geq$ 24 FPS and resolution $\geq$ 1080p. We assume that the speaker’s face is visible in recordings at a maximum viewing angle and distance of 60°$\degree$ and 3 m. Videos can be saved using common codecs (e.g., H.265, MPEG-4) and post-processed with compression, transcoding, and filter application. We assume cameras remain still throughout the speech but discuss reasonable solutions to avoid this requirement in  §12.
Video verifiers are individuals or media platforms (e.g., Facebook, YouTube, X) who seek to confirm the integrity of a video by providing it to the verification module.
Attackers disseminate falsified videos claiming to portray the speech event. We assume they possess white-box knowledge of Spotlight as well as significant computational power, and carry out attacks after a speech has taken place in an attempt to spread false media or undermine Spotlight verification. We do not consider environment-level attacks at the speech (including injection of interfering light, as discussed in §12) or tampering of Spotlight’s hardware. Attackers can perform any combination of the following:
@@
    Access all Spotlight algorithms and models, including weights.
@@
    Incorporate valid signatures from other Spotlight-protected videos into generated fake videos, i.e., replay attack.
@@
Real-world limitations   Spotlight is but one technical approach to achieving physical signatures and is not without limitations. (1) It requires deployment of the Spotlight core unit at speeches. (2) Because physical signatures are based upon a set of extracted event features, they do not address falsification of features outside this set. In this work, we develop a prototype leveraging visual features of speaker identity and lip and face motion. The prototype thus does not protect against falsifications of facial attributes (e.g., makeup) or non-facial elements (e.g, clothing, surrounding environment). Spotlight’s LSH framework is highly flexibile, supporting rich feature sets. Nonetheless, an operator must choose such features based on anticipated attacks. (3) Videos must contain the complete optical signature projection region to enable verification; as with digital signatures, videos lacking intact optical signatures are viewed as untrustworthy. Signature inclusion can be made highly likely by using a projection region close to the speaker’s face or even configuring a small but visible projection border as a cue for filmers.
While we believe that several of the aforementioned limitations can be addressed via further research (§12), we ultimately view Spotlight not as a panacea to fake speech videos, but rather a complement to passive detectors and digital signature methods. In particular, digital approaches achieve provable security by requiring consistent user and/or recording-device cooperation to bind pixel values to signatures. Spotlight and physical signatures more broadly trade provable security for flexibility, scalability, and a shift of protection agency (and efforts) from audience to speaker.
@@ 3.3. Design Challenges
Compact and pose-invariant video digest   The video digest must be extracted in real-time and capture the speaker’s identity and facial motion to protect against their falsification. Furthermore, it must be pose-invariant, so that any recording can be verified. Unfortunately, the optical channel offers limited bandwidth, constraining the digest size. This is because a receiver’s sampling rate must be at least double the modulation frequency (i.e., the Nyquist rate) for data to be decodable. The standard frame rate for recording of live events is only 30 frames per second (FPS) (framerates, ). Thus, achieving embedding bandwidths of even hundreds of bits per second (bps) is challenging. Existing speaker analysis methods output large representations (e.g., hundred-dimension embeddings for single images (Wiles18a, ) or seconds of audio (wang2018attention, )), well-exceeding this limit.
@@
We next present the design of Spotlight’s three modules, which jointly address the above challenges.
(a) Generating a signature for a window of video frames.
(b) Features from real vs. falsified videos.
Figure 4. Assembly of Spotlight signatures. (a) Feature vectors hashed via LSH underpin the compact, cryptographically-secured signature. (b) A FaceMesh signal from three speech videos: real videos from two cameras at different yaws, and a reenactment deepfake changing four speaker words. The Pearson correlation (PC) between videos’ signals reflects the similarity of their speech content.
@@ 4. Speech Video Signature Creation
We propose a feature-agnostic framework for creating compact, pose-invariant speech video signatures in real-time (Figure 4(a)). We build off existing computer vision tools to extract semantically-meaningful visual features crafted to address the attacks outlined in §2.1. Then, we use a technique known as locality-sensitive hashing (LSH) to compress these high-dimensional features to hundreds of bits (within the embedding data capacity) while preserving their verification functionality. Unlike cryptographic hash functions, which output highly different hashes for inputs with minor differences, LSH maps similar inputs to similar hashes. This enables Spotlight to validate legitimate videos despite minor feature differences inevitably arising from recording condition and feature extraction variance. While LSH is used in many domains, to our knowledge we are the first to use LSH-based features for verification.
4.1. LSH-Based Digest Framework
@@
LSH is a technique for reducing data dimensionality while preserving approximate distances between data points. An LSH scheme consists of a function $H:\mathbb{R}^{n}\mapsto\{0,1\}^{k}$ such that $D(H(\vec{u}),H(\vec{v}))$ estimates $sim(\vec{u},\vec{v})$. $sim$ is a similarity metric defined for $\vec{u},\vec{v}$, $D$ is the Hamming distance, and $k$ is a configurable hash size. An LSH scheme does not preserve the exact similarity of inputs but rather provides a probabilistic guarantee that similar inputs are mapped to similar hashes; using a larger hash size $k$ increases this probability.
@@
We use the cosine similarity LSH scheme (charikar2002similarity, ), denoted $H_{cos}$. $H_{cos}$ outputs $k$-bit hashes such that $D(H_{cos}(\vec{u}),H_{cos}(\vec{v}))$ estimates $\Theta(\vec{u},\vec{v})$. Here, $\Theta(\vec{u},\vec{v})$ is the angle between $\vec{u}$ and $\vec{v}$.
@@
The cosine similarity LSH scheme is appealing for our use case for two reasons. (1) While larger hash sizes always improve the hash’s accuracy in estimating cosine similarity, this relationship is independent of the dimensionality of input vectors. This differs from Principle Component Analysis, where information loss is proportional to dimensionality reduction. We derive an equation for the effect of hash size on hashed features’ performance (Theorem 4) to confirm this. In the context of our application, this means that our initial feature vectors can be arbitrarily high-dimensional, so long as they capture the similarity of speech videos via cosine similarity. (2) $H_{cos}$ is suitable for estimating the Pearson correlation, a popular measure of similarity in time series data. This is because the Pearson correlation of two time series is equivalent to their cosine similarity after zero-meaning. This property is key to computing our dynamic features, which must capture temporal speech characteristics.
@@
Equipped with $H_{cos}$, we can separately address the challenges of digest robustness and constraint size. Next, we describe our high-dimensional visual feature vectors, which are hashed to serve as verification data. We formally analyze this hashing method in A.5
4.2. Semantically-Meaningful Video Digests
@@
To address falsifications of speaker identity, lip and face motion, we extract two visual feature vectors: a biometric-based identity feature vector and a temporal dynamic feature vector. The LSH framework, however, supports varied features, as discussed in §12.
@@
Identity feature vector   The identity feature vector is used to verify the speaker identity in a published video to protect against identity swap falsifications. Neural network face embedding models are the gold-standard for extracting visual identity information. They map face images to vectors in a high-dimensional embedding space, where distance corresponds to face similarity. Conveniently, state-of-the-art face embedding models utilize cosine similarity as their distance metric. We employ a pre-trained ArcFace (deng2019arcface, ) model (arcface_impl, ), which outputs a 512-dimensional vector. We pass ArcFace crops of the face obtained from a pre-trained face detector (ultra_light_face_detector, ).
@@
Dynamic feature vector   The dynamic feature vector protects against falsifications of delivered content by ensuring that a speaker’s face and lip motion have not been modified. We use MediaPipe FaceMesh (mediapipe, ), a model for real-time face image analysis, to distill a window of video frames into a signal capturing both coarse and fine-grained spatio-temporal visual characteristics of the speaker. As shown in Figure 4(b), the similarity of two speech videos can be quantified as the Pearson correlation of their corresponding signals. We find that these simple signals strongly protect against varied falsifications (§9.1). They also are more compact yet comparable in robustness to other speech features, as discussed in §12.
@@
Thus, given a window of $n$ video frames, we run FaceMesh on each frame to obtain its 52 blendshape scores – pose-invariant coefficients representing facial expressions – and 478 facial landmarks which we align to a canonical view for pose-invariance. We find that FaceMesh produces accurate output for frames captured up to $60\degree$ off-axis from the speaker (§9.1). We concatenate the values of 11 blendshapes and 5 distances between landmarks around the lips into separate $n$-sample signals, which together capture global facial motion and nuanced lip motion. We identify this set of features as optimal via forward sequential feature selection (FERRI1994403, ) on all blendshape and distance signals, using a comprehensive multi-camera dataset (§9.1). Finally, we smooth and standard normalize all signals and concatenate them into our $16n$-dimensional feature vector.
@@ 4.3. MAC Generation and Key Management
Our digest consists of both feature hashes, a window number, a core unit identifier, and a creation date (Figure 4(a)). The Spotlight core unit generates a HMAC-SHA1 MAC for each digest using its secret key, ensuring the integrity and authenticity of embedded data. We refer to a digest and its MAC as a speech video signature.
We secure our data via MACs as opposed to public key encryption because public key schemes produce ciphtertexts exceeding our embedding bandwidth. [^1] For equivalent authenticity and integrity guarantees, RSA-based digital signatures are over 10 times larger than HMAC-SHA1 MACs. We discuss approaches to increasing bandwidth to facilitate public key encryption in §12.
To create and validate MACs, Spotlight must establish a secret key shared by core units and the verification service. This can be done using Diffie-Hellman key exchange (diffie2022new, ), a scheme enabling two parties to generate a shared secret key over an insecure channel using their public keys. Spotlight may employ Diffie-Hellman in one of two ways. In the first, it may require each core unit owner to use his own public key to participate in key exchange with the cloud-based verification service (§3.1). During this process, Spotlight can authenticate the owner’s public key via a digital certificate (cert_pki, ). This creates a unique secret key for each core unit, securely associated with a unit’s identifier and owner. In the second approach, Spotlight may maintain the same key across all core units, refreshing as needed. Our prototype assumes a secret key has already been initialized in one of these manners, since the involved key exchange and certificate technologies are well-established.
@@ 5. Optical Signature Embedding
After obtaining the signature for a window, the optical signature embedding module projects light encoding the signature data into the scene. Prior works most relevant to this task explore light-based (rollinglight, ; cotting2004embedding, ) or screen-camera (disco, ; lightning_markers, ; unseencode, ; imperceptible_screen_ar_markers, ; tera, ; vrcode, ; chromacode, ; deepqrcoderestoration, ; inframe, ; inframe++, ; theinvisibleqr, ; mobiscan, ; vlctpvm, ; lrsnetinvisibleqr, ; kaleido, ; wang_deepopticalwatermarking, ; kazutake_opticalwatermarking3, ; hilight, ) communication. They achieve imperceptibility at the scene while maximizing the visibility of modulated light in captured frames for real-time decoding. Such methods are inapplicable in our case, wherein we seek imperceptibility both live and in-video and decoding is performed downstream on videos rather than at capture. Further, several of these works require cameras to operate in rolling-shutter mode (disco, ; rollinglight, ).
To address these issues, we propose three design elements, illustrated in Figure 5. (1) We modulate light spatially and temporally to boost embedding bandwidth. The temporal modulation operates at low frequencies (e.g., 3-6 Hz) to accommodate commodity camera’s framerates and all shutter modes. We leverage an amplitude-modulating spatial light modulator (SLM) – an optical device that controls the intensity of emitted light in both space and time – to introduce small amounts of carefully-crafted light onto a planar surface in the immediate vicinity of the speaker. Our design maintains imperceptibility both live and in videos by exploiting the human visual system’s low sensitivity to small fluctuations in light intensity occurring in small regions (colordiscriminationvsize, ) and for short durations (colordiscriminationvtime, ). (2) We apply concatenated error correction coding to the signature data to ensure its reliable recovery from videos and enhance its resilience against video post-processing techniques. (3) We design an adaptive embedding mechanism which continually tunes the emitted light to adapt to environmental changes and balance embedding imperceptibility and robustness.

### v2

@@ 3. Preliminaries
(a) Generating a signature for a window of video frames.
(b) Features from real vs. falsified videos.
Figure 4. Assembly of VeriLight signatures. (a) The signature consists of feature vectors hashed via LSH, plus a MAC. (b) A FaceMesh signal from three speech videos: two real videos from cameras at different yaws, and a reenactment deepfake changing four speech words. The Pearson correlation (PC) between videos’ signals reflects the similarity of their speech content.
VeriLight is a physical and proactive approach to protecting live speech videos from visual falsification of speaker identity and lip and face movements, which reflect delivered content. To match the visual nature of these falsifications, we embed physical signatures with light – the sensing modality of vision. In this section, we describe VeriLight’s design and threat model, and then discuss the technical challenges it overcomes.
@@ 3.1. System Overview
Figure 3 overviews the VeriLight design, comprising three modules. The VeriLight core unit, deployed at the site of a live speech, consists of a camera observing the event and a spatial light modulator positioned to project light onto any approximately planar surface in the immediate vicinity of the speaker (e.g., a small portion of a podium, wall, backdrop, or curtain). Notably, each of the speech environments portrayed in recent high-profile falsification incidents (nan, 2019; Wade, 2023; Xander Elliards, 2024; bid, 2023; McCarthy, 2024; har, 2024b, a) possess such a surface. We believe it is rare to find a high-profile speech event where this is not the case.
For each window of camera video frames, the signature creation module running on the core unit extracts a descriptor from frames. The descriptor contains semantically-meaningful features capturing speaker identity and lip and face motion, and additional provenance metadata such as a window timestamp. The core unit then generates a message authentication code (MAC) for the descriptor with its secret key. A descriptor and its MAC comprise a signature. The optical signature embedding module encodes the signature data as modulated light projected into the scene in the subsequent window of time. This naturally embeds signatures into real recordings.
To verify a published video of a speech, the verification module extracts optical signatures from video frames and validates the MACs to confirm descriptors’ integrity. It then examines whether the semantically-meaningful features recovered from the descriptor match those of the portrayed speech. To ensure that the secret key used to validate MACs is secured, we envision the verification module as a secure cloud service queried by users and media platforms. As such, secret keys will never be exposed to third-party devices. We elaborate on VeriLight’s use of secret keys and MACs in §4.3.
@@ 3.2. Threat Model
[…]
Video verifiers are individuals or media platforms who seek to confirm a video’s integrity by passing it to the verification module.
Attackers disseminate falsified videos claiming to portray the speech event. We assume they possess white-box knowledge of VeriLight and significant computational power, and carry out attacks after a speech has taken place in an attempt to spread fake videos or undermine VeriLight verification. We counter injection of interfering light at the speech site (§10.2), but further on-site attacks, such as tampering of VeriLight’s hardware, are out of scope. Concretely, attackers can perform any combination of the following:
@@
    Access all VeriLight algorithms and models, including weights.
@@
    Incorporate valid signatures from other VeriLight-protected videos into generated fake videos, i.e., replay attack.
@@
Real-world limitations   VeriLight is but one technical approach to achieving physical signatures and is not without limitations. (1) It requires deployment of the VeriLight core unit at speeches. (2) Because physical signatures are based upon a set of extracted event features, they do not address falsification of features outside this set. We develop a prototype leveraging visual features of speaker identity and lip and face motion. The prototype thus does not protect against falsifications of facial attributes (e.g., makeup) or non-facial elements (e.g., clothing, background). VeriLight’s flexible LSH framework supports rich feature sets, but these features must be preselected based on anticipated attacks. (3) Videos must contain the full optical signature projection region to enable verification; as with digital signatures, videos lacking intact optical signatures are viewed as untrustworthy. Using a projection region close to the speaker’s face or even configuring a small but visible projection border as a cue for filmers can help ensure signature inclusion.
While we believe that several of the aforementioned limitations can be addressed via further research (§12), we ultimately view VeriLight not as a panacea to fake speech videos, but rather a complement to passive detectors and digital signature methods. In particular, digital signature methods achieve provable security by requiring consistent user and recording-device cooperation to bind pixel values to credentials. VeriLight and physical signatures more broadly trade provable security for flexibility, scalability, and a shift of protection agency (and efforts) from audience to speaker.
@@ 3.3. Design Challenges
Compact and pose-invariant video descriptor   Video descriptors must be extracted in real-time and capture the speaker’s identity and lip and facial motion. Furthermore, they must be pose-invariant, so that any recording can be verified. Unfortunately, the optical embedding channel offers limited bandwidth, constraining the descriptor size. This is because a receiver’s sampling rate must be at least double the modulation frequency (i.e., the Nyquist rate) for data to be decodable. Commodity cameras typically operate at only 24-30 frames per second (FPS). Thus, achieving embedding bandwidths of even hundreds of bits per second (bps) is challenging. Existing speaker analysis methods output large features (e.g., hundred-dimension embeddings for single images (Wiles et al., 2018) or seconds of audio (Wang et al., 2018a)), significantly exceeding this limit.
@@
We next present the design of VeriLight’s three modules, which jointly address the above challenges.
@@ 4. Speech Video Signature Creation
We propose a feature-agnostic framework for creating compact, pose-invariant speech video signatures in real-time (Figure 4(a)). We build off existing computer vision tools to extract semantically-meaningful visual features crafted to address the attacks outlined in §2.1. Then, we use a technique known as locality-sensitive hashing (LSH) to compress the high-dimensional features to hundreds of bits (within the embedding data capacity) while preserving their verification functionality. Unlike cryptographic hash functions, which output highly different hashes for inputs with minor differences, LSH maps similar inputs to similar hashes. This enables VeriLight to validate legitimate videos despite minor feature differences inevitably arising from recording condition and feature extraction variance. While prior works in video fingerprinting similarly apply LSH (Yang et al., 2004; Khelifi and Bouridane, 2017), our framework is unique in producing features that are simultaneously real-time, pose-invariant, and speech-specific.
4.1. LSH-Based Descriptor Framework
@@
LSH is a technique for reducing data dimensionality while preserving approximate distances between data points. An LSH scheme consists of a function $H:\mathbb{R}^{n}\mapsto\{0,1\}^{k}$ such that $D(H(\vec{u}),H(\vec{v}))$ estimates $sim(\vec{u},\vec{v})$. $D$ is the Hamming distance, simsim is a similarity metric defined for $\vec{u},\vec{v}$, and $k$ is a configurable hash size. An LSH scheme does not preserve the exact similarity of inputs but rather provides a probabilistic guarantee that similar inputs are mapped to similar hashes; using a larger hash size $k$ increases this probability.
@@
We use the cosine similarity LSH scheme (Charikar, 2002), denoted $H_{cos}$. $H_{cos}$ outputs $k$-bit hashes such that $D(H_{cos}(\vec{u}),H_{cos}(\vec{v}))$ estimates $\Theta(\vec{u},\vec{v})$. Here, $\Theta(\vec{u},\vec{v})$ is the angle between $\vec{u}$ and $\vec{v}$.
@@
The cosine similarity LSH scheme is appealing for our application for two reasons. (1) While larger hash sizes always improve the hash’s accuracy in estimating cosine similarity, this relationship is independent of the dimensionality of input vectors. This differs from Principle Component Analysis, where information loss is proportional to dimensionality reduction. In our use case, this means that our initial feature vectors can be arbitrarily high-dimensional, so long as their similarity is measured by cosine similarity. We confirm this by deriving a closed-form equation for the effect of hash size on our features’ performance (Theorem 4). (2) $H_{cos}$ can estimate Pearson correlation, a popular measure of time series similarity. This is because the Pearson correlation of two time series is equivalent to their cosine similarity after zero-meaning. This property is key to computing our dynamic features, which must capture temporal speech characteristics.
@@
Equipped with $H_{cos}$, we can separately address the challenges of descriptor robustness and size constraint. Below, we describe our high-dimensional visual feature vectors, which are hashed to serve as the descriptor verification data.
4.2. Semantically-Meaningful Video Descriptors
@@
To address falsifications of speaker identity and lip and face motion, we extract two visual feature vectors: a biometric-based identity feature vector and a temporal dynamic feature vector. The LSH framework, however, supports varied features, as discussed in §12.
@@
Identity feature vector   The identity feature vector is used to verify the speaker identity in a published video to protect against identity swap falsifications. Neural network face embedding models are the gold-standard for extracting visual identity information. They map face images to vectors in a high-dimensional embedding space, where distance corresponds to face similarity. Conveniently, state-of-the-art face embedding models utilize cosine similarity as their distance metric. We employ a pre-trained ArcFace (Deng et al., 2019) model (arc, 2024), which outputs a 512-dimensional vector. We pass ArcFace crops of the face obtained from a pre-trained face detector (ult, 2022).
@@
Dynamic feature vector   The dynamic feature vector protects against falsifications of delivered content by ensuring that a speaker’s face and lip motion have not been modified. We use MediaPipe FaceMesh (med, 2024), a model for real-time face image analysis, to distill a window of video frames into a signal capturing spatio-temporal visual characteristics of the speaker. As shown in Figure 4(b), the similarity of two speech videos can be quantified as the Pearson correlation of their corresponding signals. We find that these simple signals strongly protect against varied falsifications (§9.1). They also are more compact yet comparable in robustness to other speech features, as discussed in §12.
@@
Thus, given a window of $n$ video frames, we run FaceMesh on each frame to obtain its 52 blendshape scores – pose-invariant coefficients representing facial expressions – and 478 facial landmarks, which we align to a canonical view for pose-invariance. We find that FaceMesh produces accurate output for frames captured up to $60\degree$ off-axis from the speaker (§9.1). We concatenate the values of 5 distances between landmarks around the lips and 11 blendshapes into separate $n$-sample signals, which together capture global facial motion and fine-grained lip motion. We identify this set of features as optimal via forward sequential feature selection (Ferri et al., 1994) on all blendshape and distance signals, using a comprehensive multi-camera dataset (§9.1). Finally, we smooth and standardize all signals and concatenate them into our $16n$-dimensional feature vector.
@@ 4.3. MAC Generation and Key Management
Our descriptor consists of both feature hashes, a window number, a core unit identifier, and a creation date (Figure 4(a)). The core unit generates a HMAC-SHA256 MAC for each descriptor using its secret key, ensuring the integrity and authenticity of embedded data. We refer to a descriptor and its MAC as a speech video signature.
We secure our data via MACs as opposed to public key encryption because public key schemes produce ciphertexts exceeding our embedding bandwidth. [^2] For equivalent authenticity and integrity guarantees, RSA-based digital signatures are over 10 times larger than any HMAC-SHA MACs. We discuss approaches to increasing bandwidth to facilitate public key encryption in §12.
To create and validate MACs, VeriLight must establish a secret key shared by core units and the verification service. This can be done using Diffie-Hellman key exchange (Diffie and Hellman, 1976), a scheme enabling two parties to generate a shared secret key over an insecure channel using their own public-private key pairs. VeriLight may employ Diffie-Hellman in one of two ways. In the first, it may require each core unit owner to use their own public-private key pair to participate in secret key exchange with the cloud-based verification service (§3.1). During this process, VeriLight can additionally authenticate the owner by validating their public key’s digital certificate (D. Cooper, 2008). This process creates a unique secret key for each core unit, securely associated with a unit’s identifier and owner. In the second approach, VeriLight may maintain the same key across all core units, refreshing as needed. Our prototype assumes a secret key has already been initialized in one of these manners, since the involved key exchange and certificate technologies are well-established.
@@ 5. Optical Signature Embedding
After obtaining the signature for a window, the optical signature embedding module projects light encoding the signature data into the scene. Prior works most relevant to this task explore light-based (Lee et al., 2015; Cotting et al., 2004) or screen-camera (Jo et al., 2016; Cui et al., 2019; Fang et al., 2022; Woo et al., 2012; Zhang et al., 2018b; Wang et al., 2015; Unno and Uehira, 2020; Li et al., 2015) communication. They achieve imperceptibility at the scene while maximizing the visibility of modulated light in video for real-time decoding. These methods are inapplicable in our case, wherein we seek imperceptibility both live and in video, and decoding is performed downstream on videos rather than at capture time. Further, several are only compatible with rolling-shutter cameras (Jo et al., 2016; Lee et al., 2015; Woo et al., 2012).
To address these issues, we propose three design elements, illustrated in Figure 5. (1) We modulate light spatially and temporally to boost embedding bandwidth. The temporal modulation operates at low frequencies (e.g., 3-6 Hz) to accommodate commodity cameras frame rates and all shutter modes. We leverage an amplitude-modulating spatial light modulator (SLM) – an optical device that controls the intensity of emitted light in both space and time – to introduce small amounts of carefully-crafted light onto a planar surface in the immediate vicinity of the speaker. Our design maintains imperceptibility both live and in videos by exploiting the human visual system’s low sensitivity to small fluctuations in light intensity occurring in small regions (Szafir, 2017) and for short durations (Siegel, 1965). (2) We apply concatenated error correction coding to the signature data to ensure its reliable recovery from videos and enhance its resilience against video post-processing. (3) We develop an adaptive embedding algorithm which continually tunes the emitted light to respond to environmental changes and balance embedding imperceptibility and robustness.